# Multi-modal Evidential Fusion Network for Trusted PET/CT Tumor Segmentation

### v1

@@ 3.2 Cross-modality Feature Learning (CFL)
As illustrated in the top and bottom parts of Fig. 1, our proposed MEFN encompasses two CFL branches with a structure mirrors each other. Taking the CT branch as an example, given the input , the modality encoder extracts high-level feature representations of the CT image. Subsequently, the modality decoder transforms these high-dimensional features into modality-specific features and generates a synthetic PET image . Concurrently, the tumor decoder decouples tumor features from the high-dimensional features of the CT image to produce a synthetic tumor mask image . Notably, different to the segmentation task, here represents the generated pseudo tumor mask image, rather than the probability prediction map of the tumor region. The task of modality discriminator is to distinguish between and . To extract general tumor features across different modalities, a shared tumor discriminator is designed to differentiate between real tumor mask and synthetic tumor mask . The architecture of all encoders, decoders and discriminator networks adheres to the framework established by Isola et al. (2017). To further make the network focusing more on tumor features, we introduce the Multi-scale Tumor Guided attention (MTGA) module. MTGA is incorporated at every upsampling stage of , where tumor features decoupled by are sent to Global Average Pooling (GAP) and a Multi-Layer Perceptron (MLP) in MTGA module of (See Fig. 2(a)) to learn channel attention. This channel attention is then employed to weight the modality features extracted by upon element-wise multiplication (). The detailed description of process flow of MTGA is depicted in Fig. 2(a). Let represent the input, the operations of Cross-modality Feature Learning module can be mathematically formulated as follows:
where , and represent the output of the -th layer of , the output of the -th layer of and the sigmoid function, respectively. represents transposed convolution, represents the generated fake image and represents the generated tumor mask image. The entire CFL adopts adversarial training under the constraint of two adversarial losses for generating fake PET (or CT) images and for generating fake tumor mask images, respectively. Taking the CT branch as an example, the above mentioned two losses denoted as and are defined as:
where and denote - and - encoder-decoder combinations, respectively. represents the real tumor mask. Considering the difference between the real target modality and the synthetic target modality, L1 loss is used to ensure that the generated target modality is close to the real target modality, thereby reducing blurriness. The L1 loss for CT branch can be defined as,
@@
where is the hyperparameters used to balance the losses in Eq. (5). In this paper, is set as 100.
@@ 3.3 Segmentation Backbone and Multimodal Feature-levle Fusion
For convenient of parameter sharing, the network architecture of the segmentation backbone is strikingly similar to the - structure of the CFL, with the only modification being the final output layer. The CFL employs an output head denoted as (+), while the segmentation backbone utilizes an output head denoted as (+). In pursuit of facilitating multi-modal feature fusion, we propose a novel fusion network. This network introduces a specifically designed fusion branch that leverages single-modality prediction outcomes and the modality’s inherent uncertainties (details of which are delineated in section 3.4) to guide the feature fusion process. This approach not only allows for the seamless transfer of features learned from the CFL but also enables the acquisition of potent fusion features, which is instrumental for delineating the tumor regions required for segmentation.
@@
To fuse effectively features from each modality, we introduce a novel Dual-attention Feature Calibrating (DFC) module as depicted in Fig. 2(b) which employs distinct fusion strategies for the encoder and decoder features of the two mono-modality branches. For encoder features, motivated by Vaswani et al. (2017), DFC utilizes Cross Modality Attention (CMA) as illustrated in the lower left part of Fig. 2(b) which first merges features and from different layers of the two mono-modality branches using query, key and value vectors and then recovers the number of channels of the merged features through a convolution to obtain the final fused feature . Here, the query, key and value vectors are all derived from convolutions. Upon adaptive adjustments of these vectors, encoder features from each modality are selectively integrated, thereby reducing redundancy and improving fusion efficiency. For decoder features, DFC adopts an Uncertainty Calibrator (UC) as shown in the lower right part of Fig. 2(b). The UC employs the predictive uncertainties and from the two mono-modality segmentation branches as calibrators for the decoder features, thereby refining the extracted semantic information. Specifically, for decoder features and of different layers, and are first down-sampled to match the dimensions of and whose results are then multiplied in element-wise manner by convolutional results of and , respectively. The resultant ones are added residually to and respectively to alleviate degradation problem. Upon concatenating these two features in channel-wise manner and performing a convolution, the fused feature can be derived. In scenarios where the modalities exhibit areas of high uncertainty (typically at boundaries and conflicting features), UC directs the network’s focus towards these areas of high uncertainty, thereby improve the network’s precision in delineating boundary regions and reducing conflicts between modalities. Ultimately, and are connected in a manner similar to the skip connections in UNet and then fed into the multi-modal fusion branch to learn more potent fused features, culminating in the generation of a fused segmentation probability map and fused uncertainty for the tumor region.
@@
Similarly, for the segmentation results of PET and the fused segmentation result , the segmentation losses and are derived in the same manner.
3.4 Uncertainty Quantification and Dempster-Shafer Theory Based Trusted Fusion
Guided by Subjective Logic (Jsang, 2018) and based on the evidence gathered within the data, we may deduce an overall uncertainty (uncertainty masses) as well as probabilities for different categories (belief masses). Here, evidence denotes indicators collected from the input that support categorization, intimately linked to the density parameters of the Dirichlet distribution as discussed in section 3.3. Based on the evidence vector output from the network, we can ascertain the parameters of the Dirichlet distribution. Specifically, for a C-class problem, the subjective logic endeavors to allocate a belief mass to each category label and an overall uncertainty mass to the framework as a whole. Hence, for any modality , the belief mass values and the overall uncertainty are non-negative and are cumulatively equating to one:
@@ 3.4 Uncertainty Quantification and Dempster-Shafer Theory Based Trusted Fusion
Following the aforementioned fusion rules, we can obtain the estimated multi-modal joint evidence and the corresponding parameters of the joint Dirichlet distribution as well as the final prediction of the multimodal trusted fusion and the overall uncertainty . For deep supervision of network training, the final segmentation result from decision-level fusion can be obtained in a similar manner to the previously defined Eq. (14) for the loss .
@@ 4.1 Datasets and perprocessing
The proposed segmentation framework is evaluated on two public PET/CT datasets, with one originating from the 25th MICCAI Challenge on publicly available whole-body PET/CT dataset (AutoPET) (Gatidis et al., 2022) and the other from the 23rd MICCAI Challenge on head and neck tumors segmentation in PET/CT (Hecktor) (Oreiller et al., 2022).
@@
Given data from multi-centers, we normalize the CT and PET images. The CT volumes are clipped to [-1024, 1024] Hounsfield units and then mapped to [-1, 1]. We normalize the PET using the Z-score. As the segmentation is performed in 2D, we sliced the 3D datasets. Similar to the work proposed by Fu et al. (Fu et al., 2021), slices with no tumor region presented in the ground truth are excluded from the dataset in our experiments, after which we resize them to 256 256 and then add 25 negative samples to each dataset for data balance. In total, our experiments comprise 32,158 pairs of PET/CT slices from the Hecktor dataset and 29,065 pairs of PET/CT slices from the AutoPET dataset. We conducte 5-fold cross-validation, where each dataset is randomly partitioned into 5 mutually exclusive subsets based on different patients, so as to ensure that different subsets do not contain slices of the same patients. The algorithm is then run five rounds, in each of which one of the five subsets is used as the test set, one is used as the validation set while the left three are employed as the training set. Finally, the average results of five experiments on the test sets are used to evaluate the segmentation performance of the method.
@@ 4.2 Implementation details and metrics
The proposed approach is implemented using Python and the PyTorch (Paszke et al., 2019) library on a PC equipped with an NVIDIA GTX 3090 GPU (24 GB memory). The training is conducted using the Adam optimizer with a momentum of 0.99. Based on empirical observations, the learning rate is set to and the number of iterations is set to 100. By training the model with different hyperparameters and evaluating it on the validation set, the hyperparameter in Eq. (5) is manually adjusted to 100.0.
The segmentation model’s performance is evaluated in terms of the Dice similarity coefficient (DSC) in Eq. (21), Jaccard in Eq. (22), 95 Hausdorff Distance (HD95) in Eq. (23), sensitivity (Sens) in Eq. (24) and precision (Pre) in Eq. (25),
where represents the predicted output of the network and represents the ground truth labels, is the euclidean distance between points and , and are supremum and infimum, (true positives) represents the number of positive samples correctly classified as positive, (false positives) is the number of negative samples incorrectly classified as positive and (false negatives) represents the number of positive samples incorrectly classified as negative.
The Dice coefficient evaluates the similarity between the predicted labels and the ground truth by calculating the ratio of the overlapping regions. A Dice value closer to 1 indicates a higher similarity between the model prediction and the segmentation ground truth, indicating a better segmentation performance. Similarly, Jaccard is also used to measure the similarity and dissimilarity between the model output and the segmentation ground truth. A higher Jaccard value indicates better segmentation performance.
The Hausdorff distance is a measure used to quantify the distance between two sets, which computes the maximum shortest distance from any point in one set to the other set. The Euclidean distance is typically employed as the distance metric. The 95% Hausdorff Distance (95HD) refers to the maximum distance among the sorted Hausdorff distances between the point pairs of the predicted labels and the ground truth, considering only the top 95% of these distances. 95HD is commonly used to evaluate the local accuracy of predicted labels, where a smaller Hausdorff distance value indicates better segmentation results. Sensitivity measures the proportion of positives correctly identified and Precision measures the proportion of true positive pixels in the predictions.
@@
| AutoPET | CML (Zhang et al., 2021) | 78.88 27.18 | 71.25 27.83 | 3.62 12.26 | 84.00 21.29 | 83.73 25.41 |
| MMEF (Huang et al., 2021a) | 78.70 27.81 | 71.24 28.26 | 3.91 9.03 | 85.46 21.25 | 82.27 25.90 | |
| ADNet (Peng and Sun, 2023) | 80.30 28.15 | 73.72 28.77 | 3.46 8.61 | 85.68 22.07 | 85.26 25.00 | |
| AATSN (Ahmad et al., 2023) | 69.28 32.94 | 61.10 32.76 | 7.29 14.76 | 75.11 32.77 | 80.12 26.14 | |
| MFCNet (Wang et al., 2023) | 79.35 21.63 | 70.54 27.16 | 4.08 9.26 | 84.62 22.86 | 84.41 26.18 | |
| CSCA (Shu et al., 2024) | 78.71 27.52 | 71.17 28.26 | 4.09 9.48 | 88.52 21.72 | 78.76 26.15 | |
@@
| Hecktor | CML (Zhang et al., 2021) | 80.12 25.32 | 63.27 27.59 | 4.26 7.21 | 76.73 33.74 | 86.32 19.61 |
| MMEF (Huang et al., 2021a) | 78.11 23.95 | 61.11 28.63 | 4.16 7.02 | 69.09 32.14 | 86.80 17.48 | |
| ADNet (Peng and Sun, 2023) | 73.97 28.96 | 64.02 33.53 | 10.72 19.72 | 71.53 34.28 | 81.49 26.86 | |
| AATSN (Ahmad et al., 2023) | 76.73 25.21 | 65.13 26.91 | 12.24 20.85 | 78.31 27.58 | 79.43 23.13 | |
| MFCNet (Wang et al., 2023) | 74.14 27.72 | 62.96 22.41 | 6.41 7.01 | 77.20 27.13 | 76.12 28.85 | |
| CSCA (Shu et al., 2024) | 77.13 21.41 | 60.21 24.86 | 4.39 7.06 | 79.37 28.82 | 74.79 21.30 | |
@@ 4.3 Comparisons with the state-of-the-art methods
We conducted comparative segmentation experiments on both the AutoPET and Hecktor datasets. Six most recent multi-modal tumor segmentation methods, namely CML (Zhang et al., 2021), MMEF (Huang et al., 2021a), ADNet (Peng and Sun, 2023), AATSN (Ahmad et al., 2023), MFCNet (Wang et al., 2023) and CSCA (Shu et al., 2024), are compared, whose results are shown in Table 1.
@@
On the AutoPET dataset, AATSN (Ahmad et al., 2023) performs the worst among others. This may be due to its poor capability in capturing the regions of interest with its lightweight attention mechanism. As compared with the AATSN, CML (Zhang et al., 2021) achieved an improved improved performance with a DSC score of 78.88. This can be attributed to the introduction of the cross-modality Feature Transition Module to reduce the domain gap between PET and CT before segmentation. MFCNet (Wang et al., 2023) further improves the DSC score to 79.35 by integrating a calibration network but sacrifices the accuracy of tumor edges as indicated by the increased HD95 compared to CML (Zhang et al., 2021) (by 0.46m
m). ADNet (Peng and Sun, 2023) achieves high DSC score and precision (80.30 and 85.26, respectively) by learning multi-modal features through channel feature separation. Upon utilizing the Dempster-Shafer Theory for feature fusion strategy, MMEF (Huang et al., 2021a) achieves the third best sensitivity of 85.46 among all methods. CSCA (Shu et al., 2024) has the highest sensitivity of 88.52 due to its channel and space compound attention. It is worth noting that although our MEFN ranks secondly in sensitivity, it outperforms all other models in terms of metrics except the sensitivity. It achieves 82.45 in DSC, 75.32 in Jaccard, 2.96mm in HD95, 86.23 in sensitivity and 88.55 in precision. Specifically, the DSC score is improved by 2.15 compared to the second-best model, ADNet.
On the Hecktor dataset, due to its multi-task nature of the dataset, feature separation struggles to differentiate tumor features from different tasks, resulting in poor performance of ADNet (Peng and Sun, 2023), although it performs well on the AutoPET dataset. AATSN (Ahmad et al., 2023) performs well due to its lightweight attention mechanism, achieving a DSC score of 76.73 and ranking the second in Jaccard (65.13). Additionally, CML (Zhang et al., 2021) achieves high DSC score and precision (80.12 and 86.32, respectively) due to its unique multi-modal feature fusion mechanism. CSCA (Shu et al., 2024) also achieves the highest sensitivity of 79.37 due to its channel and space compound attention. Notably, our MEFN outperforms other segmentation networks significantly in terms of DSC, Jaccard, HD95, sensitivity and precision, which reaches 83.35, 71.73, 3.27mm, 78.03 and 90.06, respectively. Specifically, as compared with the second-best network, our MEFN improves the DSC score and precision by 3.23 and 3.26, respectively.
@@
Besides the aforementioned quantitative comparisons among different methods, visual assessments and comparisons are also conducted between MEFN and other state-of-the-art multi-modal tumor segmentation methods. Fig. 4 gives the segmented results from different methods for three different CT sample images of patients from AutoPET dataset where blue contour gives the predicted tumor region while red contour gives the ground truth. Fig. 5 gives the segmentation results for three CT sample images of three patients from the Hecktor dataset where blue and yellow contours give the predicted regions for GTVp and GTVn tumors while red and green contours give the ground truths for these two kinds of tumors, respectively. Tumor regions marked with purple boxes are enlarged and placed on the right bottom of each sample image. Red arrows point out the mis-segmented tumor regions in each sample image. From Fig. 4, it can be seen that for small tumors throughout the body, such as in the first patient, CML (Zhang et al., 2021) and AATSN (Ahmad et al., 2023) lead to under segmented results, while MMEF (Huang et al., 2021a) and MFCNet (Wang et al., 2023) tend to merge two separate tumors into one large tumor. Although CSCA (Shu et al., 2024) can segment all tumors, it fails to delineate the tumor edges accurately. Additionally, for the second and third patients, the segmentation results of other models also differ in shape from the ground truth and exhibit many redundant edges. In contrast, our MEFN not only effectively segments all small tumors but also achieves the highest accuracy in fitting the tumor edges. It can also be seen in Fig. 5 that most methods fail to accurately distinguish tumorous and normal tissues in boundary regions, which results in redundant or missing tumor edges such as in the segmentation results of the first and second patients. Furthermore, for regions with adjacent tumors, such as in the third patient, AATSN (Ahmad et al., 2023) cannot accurately locate and segment small tumors, while other methods tend to segment two separate tumors into one larger tumor. In comparison, our method can precisely locate each tumor and capture subtle differences between tumors, yielding better segmentation results. These results indicate that our proposed MEFN achieves more accurate segmentation masks and more precise tumor edges compared to other methods.

### v2

@@ 3.2 Cross-modality Feature Learning (CFL)
As illustrated in the top and bottom parts of Fig. 1, our proposed MEFN encompasses two CFL branches with a structure mirrors each other. Taking the CT branch as an example, given the input , the modality encoder extracts high-level feature representations of the CT image. Subsequently, the modality decoder transforms these high-dimensional features into modality-specific features and generates a synthetic PET image . Concurrently, the tumor decoder decouples tumor features from the high-dimensional features of the CT image to produce a synthetic tumor mask image . Notably, different to the segmentation task, here represents the generated pseudo tumor mask image, rather than the probability prediction map of the tumor region. The task of modality discriminator is to distinguish between and . To extract general tumor features across different modalities, a shared tumor discriminator is designed to differentiate between real tumor mask and synthetic tumor mask . The architecture of all encoders, decoders and discriminator networks adheres to the framework established by Zhu et al. [32]. To further make the network focusing more on tumor features, we introduce the Tumor Guided attention (TGA) module. TGA is incorporated at every upsampling stage of , where tumor features decoupled by are sent to Global Average Pooling (GAP) and a Multi-Layer Perceptron (MLP) in TGA module of (See Fig. 2(a)) to learn channel attention. This channel attention is then employed to weight the modality features extracted by upon element-wise multiplication (). The detailed description of process flow of TGA is depicted in Fig. 2(a).
The entire CFL adopts adversarial training under the constraint of two adversarial losses for generating fake PET (or CT) images and for generating fake tumor mask images, respectively. Taking the CT branch as an example, the above mentioned two losses denoted as and are defined as:
where and denote - and - encoder-decoder combinations, respectively. , , , , represent the real CT, real PET, fake PET, real tumor mask and predicted tumor mask, respectively. Considering the difference between the real target modality and the synthetic target modality, L1 loss is used to ensure that the generated target modality is close to the real target modality, thereby reducing blurriness. The L1 loss for CT branch can be defined as,
@@
where is the hyperparameters used to balance the losses in Eq. (4).
@@ 3.3 Segmentation Backbone and Multimodal Feature-levle Fusion
For convenient of parameter sharing, the network architecture of the segmentation backbone is strikingly similar to the - structure of the CFL, with the only modification being the final output layer. The CFL employs an output head denoted as (+), while the segmentation backbone utilizes an output head denoted as (+). In pursuit of facilitating multi-modal feature fusion, we propose a novel fusion network. This network introduces a specifically designed fusion branch that leverages single-modality prediction outcomes and the modality’s inherent uncertainties (details of which are delineated in section 3.4) to guide the feature fusion stage. This approach not only allows for the seamless transfer of features learned from the CFL but also enables the acquisition of potent fusion features, which is instrumental for delineating the tumor regions required for segmentation.
@@
To fuse effectively features from each modality, we introduce a novel Dual-attention Feature Calibrating (DFC) module as depicted in Fig. 2(b) which employs distinct fusion strategies for the encoder and decoder features of the two mono-modality branches. For encoder features, motivated by Vaswani et al. [33], DFC utilizes Cross Modality Attention (CMA) as illustrated in the lower left part of Fig. 2(b) which first merges features and from different layers of the two mono-modality branches using query, key and value vectors and then recovers the number of channels of the merged features through a convolution to obtain the final fused feature . Here, the query, key and value vectors are all derived from convolutions. Upon adaptive adjustments of these vectors, encoder features from each modality are selectively integrated, thereby reducing redundancy and improving fusion efficiency. For decoder features, DFC adopts an Uncertainty Calibrator (UC) as shown in the lower right part of Fig. 2(b). The UC employs the predictive uncertainties and from the two mono-modality segmentation branches as calibrators for the decoder features, thereby refining the extracted semantic information. Specifically, for decoder features and of different layers, and are first down-sampled to match the dimensions of and whose results are then multiplied in element-wise manner by convolutional results of and , respectively. The resultant ones are added residually to and respectively to alleviate degradation problem. Upon concatenating these two features in channel-wise manner and performing a convolution, the fused feature can be derived. In scenarios where the modalities exhibit areas of high uncertainty (typically at boundaries and conflicting features), UC directs the network’s focus towards these areas of high uncertainty, thereby improve the network’s precision in delineating boundary regions and reducing conflicts between modalities. Ultimately, and are connected in a manner similar to the skip connections in UNet and then fed into the multi-modal fusion branch to learn more potent fused features, culminating in the generation of a fused segmentation probability map and fused uncertainty for the tumor region.
@@
Similarly, for the segmentation results of PET and for the fused segmentation result, the segmentation losses and are derived in the same manner.
3.4 Uncertainty Quantification and Dempster-Shafer Theory Based Trustworthy Fusion
Guided by Subjective Logic [34] and based on the evidence gathered within the data, we may deduce an overall uncertainty (uncertainty masses) as well as probabilities for different categories (belief masses). Here, evidence denotes indicators collected from the input that support categorization, intimately linked to the density parameters of the Dirichlet distribution as discussed in section 3.3. Based on the evidence vector output from the network, we can ascertain the parameters of the Dirichlet distribution. Specifically, for a C-class problem, the subjective logic endeavors to allocate a belief mass to each category label and an overall uncertainty mass to the framework as a whole. Hence, for any modality , the belief mass values and the overall uncertainty are non-negative and are cumulatively equating to one:
@@ 3.4 Uncertainty Quantification and Dempster-Shafer Theory Based Trustworthy Fusion
Following the aforementioned fusion rules, we can obtain the estimated multi-modal joint evidence and the corresponding parameters of the joint Dirichlet distribution as well as the final prediction of the multimodal Trustworthy fusion and the overall uncertainty . For deep supervision of network training, the final segmentation result from decision-level fusion can be obtained in a similar manner to the previously defined Eq. (13) for the loss .
@@ 4.1 Datasets and perprocessing
The proposed segmentation framework is evaluated using two public PET/CT datasets: the AutoPET dataset from the 25th MICCAI Challenge on whole-body PET/CT segmentation [35] and the Hecktor dataset from the 23rd MICCAI Challenge on head and neck tumor segmentation in PET/CT [36].
@@
Given data from multi-centers, we normalize the CT and PET images. The CT volumes are clipped to [-1024, 1024] Hounsfield units and then mapped to [-1, 1]. We normalize the PET using the Z-score. As the segmentation is performed in 2D, we sliced the 3D datasets. After which we resize them to 256 256 and then add 25 negative samples to each dataset for data balance. In total, our experiments comprise 32,158 pairs of PET/CT slices from the Hecktor dataset and 29,065 pairs of PET/CT slices from the AutoPET dataset. We conducte 5-fold cross-validation, where each dataset is randomly partitioned into 5 mutually exclusive subsets based on different patients, so as to ensure that different subsets do not contain slices of the same patients.
@@ 4.2 Implementation details and metrics
The proposed approach is implemented using Python and the PyTorch [37] library on a PC equipped with an NVIDIA GTX 3090 GPU (24 GB memory). The training is conducted using the Adam optimizer with a momentum of 0.99. Based on empirical observations, the learning rate is set to and the number of epoch is set to 100. The parameter in Eq. (4) is determined upon grid search on candidate values {1, 50, 100, 150, 200}. The value of 100 with the highest validation accuracy is chosen for our subsequent experiments. The segmentation model’s performance is evaluated in terms of the Dice similarity coefficient (DSC), Jaccard, 95 Hausdorff Distance (HD95), sensitivity (Sens) and precision (Pre).
@@
AutoPET | CML (2021) [38] | 78.88 27.18 | 71.25 27.83 | 3.62 12.26 | 84.00 21.29 | 83.73 25.41 |
MMEF (2021) [31] | 78.70 27.81 | 71.24 28.26 | 3.91 9.03 | 85.46 21.25 | 82.27 25.90 | |
SCL-Net (2023) [39] | 71.35 32.51 | 63.46 32.25 | 5.32 11.05 | 76.99 30.52 | 80.30 26.60 | |
AATSN (2023) [30] | 69.28 32.94 | 61.10 32.76 | 7.29 14.76 | 75.11 32.77 | 80.12 26.14 | |
MFCNet (2023) [40] | 79.35 21.63 | 70.54 27.16 | 4.08 9.26 | 84.62 22.86 | 84.41 26.18 | |
Mirror U-Net (2023) [41] | 75.80 30.08 | 68.22 30.24 | 4.32 10.10 | 81.06 27.59 | 82.83 24.03 | |
@@
Hecktor | CML (2021) [38] | 80.12 25.32 | 63.27 27.59 | 4.26 7.21 | 76.73 33.74 | 86.32 19.61 |
MMEF (2021) [31] | 78.11 23.95 | 61.11 28.63 | 4.16 7.02 | 69.09 32.14 | 86.80 17.48 | |
SCL-Net (2023) [39] | 72.42 24.03 | 62.84 27.46 | 9.80 11.82 | 71.84 34.62 | 75.98 28.63 | |
AATSN (2023) [30] | 76.73 25.21 | 65.13 26.91 | 12.24 20.85 | 78.31 27.58 | 79.43 23.13 | |
MFCNet (2023) [40] | 74.14 27.72 | 62.96 22.41 | 6.41 7.01 | 77.20 27.13 | 76.12 28.85 | |
Mirror U-Net (2023) [41] | 76.07 23.49 | 61.56 26.99 | 4.55 9.91 | 71.50 33.27 | 73.56 25.16 | |
@@ 4.3 Comparisons with the state-of-the-art methods
We conducted comparative segmentation experiments on both the AutoPET and Hecktor datasets. Six most recent multi-modal tumor segmentation methods, namely CML [38], MMEF [31], SCL-Net [39], AATSN [30], MFCNet [40] and Mirror U-Net [41], are compared, whose results are shown in Table 1.
@@
On the AutoPET dataset, CML [38] achieved an improved performance with a DSC score of 78.88 and Jaccard score of 71.25. This can be attributed to the introduction of the cross-modality Feature Transition Module to allevate the impact of domain gap between PET and CT before segmentation. MFCNet [40] further improves the DSC score to 79.35 by integrating a calibration network but sacrifices the accuracy of tumor edges as indicated by the increased HD95 compared to CML [38] (by 0.46mm). Upon utilizing the Dempster-Shafer Theory for feature fusion strategy, MMEF [31] achieves the second best sensitivity of 85.46 among all methods. It is worth noting that our MEFN outperforms all other models in terms of metrics. Specifically, the DSC score is improved by 3.10 compared to the second-best model, MFCNet [40]. On the Hecktor dataset, AATSN [30] performs well due to its lightweight attention mechanism, achieving a sensitivity score of 78.31 and ranking the second in Jaccard (65.13). Additionally, CML [38] achieves high DSC score and precision (80.12 and 86.32, respectively) due to its unique multi-modal feature fusion mechanism. Notably, our MEFN outperforms other segmentation networks. Specifically, as compared with the second-best network, our MEFN improves the DSC score by 3.23.
@@
Besides the aforementioned quantitative comparisons among different methods, visual assessments and comparisons are also conducted between MEFN and other state-of-the-art multi-modal tumor segmentation methods. Fig. 4 gives the segmented results from different methods for three different CT sample images of patients from AutoPET dataset where blue contour gives the predicted tumor region while red contour gives the ground truth. Fig. 5 gives the segmentation results for three CT sample images of three patients from the Hecktor dataset where blue and yellow contours give the predicted regions for GTVp and GTVn tumors while red and green contours give the ground truths for these two kinds of tumors, respectively. Tumor regions marked with purple boxes are enlarged and placed on the right bottom of each sample image. Red arrows point out the mis-segmented tumor regions in each sample image. From Fig. 4, it can be seen that for small tumors throughout the body, such as in the first patient, CML [38], SCL-Net [39], Mirror U-Net [41] and AATSN [30] lead to under segmented results, while MMEF [31] and MFCNet [40] tend to merge two separate tumors into one large tumor. Additionally, for the second and third patients, the segmentation results of other models also differ in shape from the ground truth and exhibit many redundant edges. In contrast, our MEFN not only effectively segments all small tumors but also achieves the highest accuracy in fitting the tumor edges.
The performance improvements of the MEFN algorithm stem from three key enhancements. First, MEFN introduces a novel modality translation stage, CFL, which aligns features across modalities and learns more robust feature representations. Unlike traditional approaches, which struggle to effectively focus on tumor regions, MEFN incorporates TGA, a targeted strategy that enhances the network’s attention to these critical areas. Second, MEFN employs an advanced attention mechanism for feature fusion via the DFC module. This approach outperforms conventional concatenation-based methods [31] and single attention mechanisms [30]. Within DFC, the MCA in the encoder and UC in the decoder leverage mutual attention and uncertainty calibration to minimize redundant and conflicting features, significantly improving segmentation efficiency. Finally, the DTF module optimizes decision-level fusion by assigning weights based on uncertainty. Results with higher uncertainty are given lower weights, ensuring more reliable segmentation outcomes.